OpenStudy (18jonea):

Recall that a quadratic equation is written in the form of ax2 + bx + c = 0. For each equation below, identify a, b, and c. b. Determine the value of the discriminant: b2 – 4ac.

OpenStudy (18jonea):



OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

x would be b

OpenStudy (michele_laino):

hint: I rewrite the first equation like below: \((+1)\cdot x^2 + (-4) \cdot x+ (-5)=0\) now, please compare such equation with the general equation: \(a \cdot x^2 + b \cdot x+ c=0\)

OpenStudy (18jonea):

what is a b and c though

OpenStudy (michele_laino):

yes! what are: \(a=...?\) \(b=...?\) \(c=...?\)

OpenStudy (18jonea):

a=1 b=4 c=5?

OpenStudy (michele_laino):

please look at the coefficients inside the parentheses

OpenStudy (18jonea):

sorry so b=-4 c=-5?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, the corresponding discriminant is: \(\Delta= b^2-4ac= (-5)^2-[4 \cdot (+1) \cdot (-5)]=...?\)

OpenStudy (michele_laino):

oops.. sorry, here is the right formula: \(\Delta= (-4)^2-[4 \cdot (+1) \cdot (-5)]=...?\)

OpenStudy (18jonea):

so -4^2- 4 x1 x -5= 36 right?

OpenStudy (michele_laino):

that's right! \(\Delta=16+20=36\)

OpenStudy (18jonea):

cause 16- (-20)=36

OpenStudy (michele_laino):

yes!

OpenStudy (18jonea):

what does the triangle stand for

OpenStudy (michele_laino):

\(\Delta\) is the greek letter \(delta\)

OpenStudy (18jonea):

do I have to find square roots or no

OpenStudy (18jonea):

cause the last part states determine whether the solution will be two real solutions, one real solution, or no real solution; two imaginary solutions.

OpenStudy (michele_laino):

I don't think, since the exercise asks for the coefficients and discriminant only. In order to answer to last part, we have to establish the sign of \(\Delta\). Now, since \(36>0\), then from the general theory we can state that there are two real solutions, and such solutions are different each from other

OpenStudy (18jonea):

ok and how would you find the solutions

OpenStudy (michele_laino):

in order to find the solutions, I apply this formula: \[\Large x = \frac{{ - b \pm \sqrt \Delta }}{{2a}} = \frac{{ - \left( { - 4} \right) \pm \sqrt {36} }}{{2 \cdot 1}} = ...?\]

OpenStudy (18jonea):

-1 and 5

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (18jonea):

ok Thank You! I have A couple more of these Problems that I am going to do and Then Could you check them for me

OpenStudy (michele_laino):

ok!

OpenStudy (18jonea):

8x^2+40x+50=0 a=8 b=40 c=50 40^2-4x8x50 1,600- 1,600=0 One solution but could you show me how i find that solution

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

2x^2 +x+28=0 a=2 b=1 c=28 1-4x2x28= -911 no real solutions

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

more precisely, we have: \(\Delta=1-56 \cdot 4=1-224=-223<0\)

OpenStudy (18jonea):

how would I find the solution in the one equation?

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

since \(\Delta<0\), in order to compute the solutions, we have to introduce the complex numbers. Complex numbers are numbers \(z\) like below: \(z=a+ib\) where \(a,\;b\) are real numbers, and \(i\) is such that \(i^2=-1\)

OpenStudy (18jonea):

I got it Thank You! Could You help with a complex number thing?

OpenStudy (michele_laino):

ok!

OpenStudy (18jonea):

Do you want me to post a new Question?

OpenStudy (michele_laino):

ok! :)